You are given the following budgeted and actual data for the Grey Company for each of the months January through June of the current year.

In December of the prior year, sales were forecasted as follows: January, 93 units; February, 88 units; March, 95 units; April, 100 units; May, 107 units; June, 115 units. In January of the current year, sales for the months February through June were reforecasted as follows: February, 83 units; March, 95 units; April, 95 units; May, 97 units; June, 110 units. In February of the current year, sales for the months March through June were reforecasted as follows: March, 90 units; April, 95 units; May, 92 units; June, 110 units. In March of the current year, sales for the months April through June were reforecasted as follows: April, 95 units; May, 87 units; June, 100 units. In April of the current year, sales for the months May and June were reforecasted as follows: May, 77 units; June, 95 units. In May of the current year, sales for June were reforecasted as 95 units.

Actual sales for the six-month period, January through June, were as follows: January, 82 units; February, 83 units; March, 94 units; April, 99 units; May, 101 units; June, 110 units.

Required:

1. Prepare a schedule of forecasted sales, on a rolling basis, for the months January through June, inclusive. (Hint: There will be only one forecasted number for January—this is the forecast done in December. For February, there will be two forecasts: one done in December and a second done in January. For June, there will be six forecasts, one done in each of the preceding six months.)

2. For each of the months March through June, determine the 3-month forecast error rate, defined as 1 minus the absolute percentage error. For example, the forecast error rate for March’s sales is found by dividing the absolute value of the forecast error for this month by the actual sales volume for the month. The forecast error for any month (e.g., March) is defined as the difference between the actual sales volume for the month and the sales volume for that month forecasted 3 months earlier (e.g., December). Also, indicate for each month whether the actual sales volume was above or below the forecasted volume generated three months earlier.
Answer #1
SCHEDULE OF FORECAST SALES
1 January February March April May June
Actual Sales 82 83 94 99 101 110
Forecast in December 93 88 95 100 107 115
Forecast in January 83 95 95 97 110
Forecast in February 90 95 92 110
Forecast in March 95 87 100
Forecast in April 77 95
Forecast in May 95
2 Three Month Forecast Error Rate
a 3 Month Forecast Error for March 1 (95-94) December Forecast= 95
b 3 Month Forecast Error for April 4 (99-95) January Forecast= 95
c 3 Month Forecast Error for May 9 (101-92) February Forecast 92
d 4 Month Forecast Error for June 10 (110-100) March Forecast 100
e Actual Sales in March 94
f Actual Sales in April 99
g Actual Sales in May 101
h Actual Sales in June 110
i=a/e Forecast Error Rate For March 1.06% Actual Sales Volume below Forecast
j=b/f Forecast Error Rate For April 4.04% Actual Sales Volume above Forecast
k=c/g Forecast Error Rate For May 8.91% Actual Sales Volume above Forecast
l=d/h Forecast Error Rate For June 9.09% Actual Sales Volume above Forecast